OpenStudy (anonymous):

convert 68.3cm^3 to cubic meters

OpenStudy (gunsodo):

Divide 68.3 by 10^6

OpenStudy (anonymous):

what 10^6 referring as?

OpenStudy (anonymous):

1 cm = 1/100 m Write 68.3cm^3 as 68.3 (1 cm)^3 Put 1 cm = 1/100 m you get 68.3cm^3 = 68.3 * (1/100 m) ^3 = 68.3* 1/1000000 m^3 1000000 = 10^6 So 68.3 cm^3 = 68.3*1/10^6 m^3

OpenStudy (gunsodo):

10^6 refers as 10 x 10 x 10 x 10 x 10 x 10 or 10 multiplies by itself for 6 times

OpenStudy (gunsodo):

The final answer is 0.0000683

OpenStudy (anonymous):

okay let me try

OpenStudy (anonymous):

okay, i got the steps. Now another question : 556mL to dm^3

OpenStudy (gunsodo):

Okay, first a dm^3 equals 1 liter. So 1 liter = 1,000 mL. And then, divide 556 by 1000. You'll get the answer.

OpenStudy (anonymous):

0.556 dm^3?

OpenStudy (gunsodo):

Yes, you're correct.

OpenStudy (anonymous):

Okay another one : 200cm^3 to nm^3

OpenStudy (anonymous):

Take 1 cm = 10^7 nm and do the steps.

OpenStudy (anonymous):

okay, i'll try

OpenStudy (anonymous):

how can i know that 1 cm is equal to 10^7 nm? Is there a way to know about it?

OpenStudy (anonymous):

It's a fact. \[ 1 nm = 10^{-9} m\]

OpenStudy (gunsodo):

nm is a prefix that means like what Mr. ShailKumar has said. n stands for nano- , so, nano- means 10^-9

OpenStudy (gunsodo):

http://en.wikipedia.org/wiki/Metric_prefix Check this out!!

OpenStudy (anonymous):

i'm checking the site

OpenStudy (anonymous):

i mean how to change the 10^-9 to 10-7?

OpenStudy (anonymous):

1 m = 10^9 nm is equivalent to 100 cm = 10^9 nm 1 cm = 10^9 nm /100 1 cm = 10^7 nm

OpenStudy (anonymous):

1m = 10^-9nm?

OpenStudy (anonymous):

No. 1m = 10^9 nm i.e 1nm = 10^-9 m

OpenStudy (anonymous):

oohh i get it

OpenStudy (anonymous):

can you help me with this : (10^7)^3

OpenStudy (anonymous):

(10^7)^3 = 10^7*3 = 10^21

OpenStudy (gunsodo):

(10^7)^3 =10^(7x3) =10^21

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2 x 10^-19?

OpenStudy (gunsodo):

Yes, you're right.

OpenStudy (anonymous):

so, 1 kg = 10^6 microgram 1 microgram = 10^-6 gram ??

OpenStudy (anonymous):

Yes, if it is 1 gram = 10^6 microgram

OpenStudy (anonymous):

can you help me with changing this : 11.3 microgram to kg? I'm still confuse

OpenStudy (anonymous):

11.3 microgram = 11.3 * 1 microgram now 1 microgram = 10^-6 gram So, 11.3 mg = 11.3 * 10^-6 gram = 11.3*10^-6 * 1gram Again 1gram = 10-3 kg So 11.3 mg = 11.3*10^-6 * 10^-3 kg i.e 11.3 mg = 11.3 * 10 ^(6+3) kg = 11.3*10^-9 kg

OpenStudy (anonymous):

Correction :i.e 11.3 mg = 11.3 * 10 ^-(6+3) kg = 11.3*10^-9 kg